# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

### v1

@@ 3.2 CDSS adopting Post Hoc XAI Methodologies on top of Intrinsic Interpretable Models
Post-hoc methodologies can be applied on top of intrinsic interpretable models to provide a further layer of explainability. In this subfield, we surveyed the works of Barda et al. (2020), Kim et al. (2022), Kumarakulasinghe et al. (2020), Kovalchuk et al. (2022), Matthiesen et al. (2021), Rajashekar et al. (2024), and Sivaraman et al. (2023). All these works employed model-agnostic methodologies.
@@
Barda et al. (2020) developed a CDSS to prevent in-hospital mortality in pediatric intensive care units (PICU). The CDSS leveraged a random forest, and SHAP values on top, to predict PICU admission. Overall, feedback from focus group sessions with clinicians indicated that the proposed solution was perceived as useful. However, the capacity to digest the information varied across clinical roles and levels of predictive modeling knowledge. In general, they preferred solutions that required less cognitive effort and could accommodate diverse user information needs. Still adopting ensemble methods, Kumarakulasinghe et al. (2020) analyzed the agreement rate between physicians and LIME feature importances for sepsis prediction. Firstly, the authors trained several models such as random forest, adaboost, among others, and applied the LIME framework on top of the best-performing one (random forest). Secondly, physicians were asked to: (1) agree with the model whether the patient presented sepsis, based on individual features, and (2) blindly rank the top 3 features the model used to make predictions without accessing LIME explanations. Finally, LIME explanations were revealed and physicians were asked to rate their satisfaction on a 5-stars Likert scale. Results indicated that physicians agreed with model predictions (87%), and that LIME explanations were equal to the physicians’ explanations in at least two features for all patients (69%). Also, physicians were satisfied with LIME predictions (78%), and 68% presented a positive attitude towards explainability in terms of trust. However, such positive attitudes should not be translated into increased trust, as they found inconsistency in LIME explanations. Similarly in the scope of sepsis prediction, Sivaraman et al. (2023) examined clinicians’ interactions with an AI-based interpretable treatment recommender. The system was modeled using XGBoost, and SHAP as feature explainer. Through a think-aloud study involving 24 intensive care clinicians, the analysis revealed that explanations improved clinicians’ perceptions of the AI’s usefulness. In particular, the study identified four distinct behavior patterns among clinicians summarised as: Ignore, Negotiate, Consider, and Rely. Overall, clinicians who found the AI recommendations most useful saw it as additional evidence alongside their assessments.
@@
In the field of medical wearable technologies, Kim et al. (2022) developed MindScope, a system based on a personalized ML system trained on an XGBoost to predict patients’ stress levels. Interviews were conducted to gather feedback on the relevancy of feature importance (SHAP values) in three different explanation settings: (1) no explanations, (2) broad explanations, and (3) detailed explanations. Overall, it was found that the explanation-based settings were more useful than no explanations. However, detailed explanations were perceived as equally useful as broad explanations, thus questioning how detailed explanations should be.
@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
Next, in the field of oncology, Sabol et al. (2020) adopted a Cumulative Fuzzy Class Membership Criterion (CFCMC) to predict colorectal cancer. They developed two segmentation systems for whole-slide histopathological images: one using a standalone CNN and the other using the CFCMC classifier on CNN-extracted features to provide explanations. Through surveys with 14 pathologists, the CFMC showed higher acceptability and trust compared to a plain CNN, pointing out that explainable classifiers improve AI usability and reliability in medical settings.
@@ 4 Challenges and Future Directions
4.1 How XAI Impacts CDSS: a Socio-Technical Perspective
@@
We have reviewed the diversity of CDSS that adopt XAI human-centered evaluations in the medical field. Given the high-risk environment, application-grounded evaluations with domain experts were the main evaluation methodology. On average, clinical perceptions of XAI-based CDSS were mostly positive. For example, Ellenrieder et al. (2023) demonstrated that providing explanations not only enhances clinicians’ CDSS learning but also reduces false learning. In particular, Neves et al. (2021) demonstrated that less experienced practitioners could benefit from XAI explanations, while Brennan et al. (2019) revealed that adding XAI aids in significant improvements in physicians’ risk assessments.
However, incorporating XAI still poses significant challenges. For instance, Barda et al. (2020) discovered that the capacity to absorb explanations varied across clinical roles and levels of AI knowledge, thus advocating solutions that demand less cognitive effort. For the same reason, Kim et al. (2022) found that detailed explanations were perceived as equally useful as broad explanations, thus questioning how deep explanations should extend. Again, Kyrimi et al. (2020) documented that too-engineered explanations led to brain clutter, preventing their full internalization.
@@
Notwithstanding, Sivaraman et al. (2023) more granularly identified that clinicians who ignored explanations were already confident in their own settled decisions, but clinicians who “negotiated” accepting at least one aspect of the recommendation were able to reach a balanced intermediate solution that combined with their own intuition. Similarly, Matthiesen et al. (2021) argued that ML and explainability did not surpass specialized and experienced clinicians’ capabilities, but at best support them during the clinical workflow, as well as Kyrimi et al. (2020), who found that, although XAI was useful, current conventional clinical practices are still perceived as more understandable. Finally, we highlight that only Hwang et al. (2022); Matthiesen et al. (2021) early gathered clinicians’ feedback before CDSS development, which may be crucial to ensure that the CDSS aligns with real-world clinical expectations, thus purportedly reducing frictions in subsequent stages. These findings emphasize the need for explanations to align with prior clinical knowledge through early stakeholder feedback and to minimize the cognitive burden on clinicians using CDSS.
@@
More in general, despite explainability methods face intrinsical issues such as lack of robustness Slack et al. (2020) and intra-method disagreement Krishna et al. (2022), issues that may be framed as independent research topics per sé, the friction from clinicians in perceiving explainability may be attributed to socio-technical gaps between human requirements and the effective utility of technological solutions, because human activity is “flexible and nuanced”, while computational mechanisms are “rigid and brittle” Ackerman (2000). In other words, a socio-technical gap refers to the discrepancy between human needs, including expectations, and the technological capabilities designed to support them, highlighting the limitations of technological systems in fully accommodating the nuances of human interactions. This general theoretical construct can be extended to the purpose of XAI explanations in CDSS, as explanation capabilities might not meet rigorous human requirements from precise clinical knowledge, given that CDSS are dramatically at high stake, as output recommendations may have a significant impact on patient lives. In other words, implementing XAI solutions in CDSS might not always bridge the socio-technical gap when compared to systems that do not adopt XAI, or systems that use basic reasoning in current clinical knowledge.
@@
4.2 A Medical Field-Agnostic, Stakeholder-Centric Iterative Approach to CDSS Development
The iterative process can be decomposed into the following phases: (1) stakeholder identification, (2) stakeholders’ requirements identification, and (3) system development and evaluation.
First, it is fundamental to identify all relevant stakeholders in each specific case. As a rule of thumb, identifiable stakeholders in the healthcare domain may be: (1) patients, (2) clinicians, (3) clinics and hospitals - including their corporate governance and management, and (4) CDSS developers - including researchers and engineers. Patients may be considered as passive stakeholders, as they do not directly interact, but are “targeted” by the system. On the contrary, clinicians are actively supported by CDSS, corporate management may decide whether the clinic/hospital should adopt AI-based solutions, evaluating the costs and benefits, and developers are responsible for the creation and maintenance of the technological workflows.
Second, we argue that a participatory AI methodology in which all stakeholders are early on involved is an effective starting point, where diverse requirements are early considered, potentially reducing frictions afterward, e.g., Hwang et al. (2022); Matthiesen et al. (2021). This entails gathering all stakeholders’ needs, setting goals, and clearly framing an evaluation strategy specific to the CDSS and to the medical field in which the CDSS would operate. To this end, we suggest developers inform and instruct non-technical stakeholders of how CDSS operate ex-ante, including carefully describing their inner workings based on machine learning model(s) and XAI explainability component(s), and possible evaluation frameworks. Here, presenting them with relevant taxonomies can help them become familiar with the material more effectively (see Section 2.2 for XAI principles and methods, and 2.3 for human-centered evaluations). Methodologies may include a pilot study for existing prototypes or a formative study if none exist. This phase should capture stakeholders’ expectations and assess the need for specialized evaluation methods in the CDSS’s medical field. Also, focus groups can be particularly useful, as they convene all stakeholders simultaneously.
Third, CDSS developers would optimally internalize stakeholders’ requirements gathered from the previous phase to develop a prototype to be quickly tested with synthetic patients’ data. The prototype should adhere to both technical backend requirements and usability requirements and be designed with a focus on future adoption, thus seamlessly integrable into medical workflows. We argue that concurrent think-aloud studies can be effective in this stage, as clinicians (the end users) can articulate their thoughts on the perception of the accuracy and explainability of the models.
In general, steps 2 and 3 should be conceived as iterative, adhering to the concept that continuous feedback stimulates CDSS improvement. This iterative process guarantees that system performance is continuously and rigorously improved. Through the incorporation of ongoing feedback, the CDSS can be fine-tuned with new information, thus refining both its predictive value and clinical usefulness. As such, this refinement process facilitates evidence-based decision-making and optimizes overall system effectiveness in clinical practice.
4.3 Conclusion
@@
In this work, we have examined the role of explainability in CDSS and the socio-technical challenges it faces. While explainability enhances clinicians’ trust, its effectiveness varies based on user expertise, cognitive load, and the depth of explanation required. The socio-technical gap remains a central challenge, as human decision-making is inherently flexible, whereas (X)AI models operate within rigid constraints. Given the high stakes in healthcare, rigorous evaluation strategies - particularly application-grounded methodologies - are necessary to ensure that CDSS align with stakeholders’ requirements. To address these challenges, we proposed an iterative framework for CDSS development to integrate XAI and HCI principles. By identifying key stakeholders, gathering their requirements through participatory methodologies, and continuously refining CDSS through iterative evaluations, we advocate for a development approach that prioritizes usability, trust, and clinical integration. Overall, rather than replacing human expertise, CDSS should serve as augmentative tools that enhance clinicians’ decision-making, ultimately fostering more effective and informed patient care.

### v2

@@ 3.2 CDSS adopting Post Hoc XAI Methodologies on top of Intrinsic Interpretable Models
Post-hoc methodologies can be applied on top of intrinsic interpretable models to provide a further layer of explainability. In this subfield, we surveyed the works of Barda et al. (2020), Kim et al. (2022), Kumar. et al. (2020), Kovalchuk et al. (2022), Matthiesen et al. (2021), Rajashekar et al. (2024), and Sivaraman et al. (2023). All these works employed model-agnostic methodologies.
@@
Barda et al. (2020) developed a CDSS to prevent in-hospital mortality in pediatric intensive care units (PICU). The CDSS leveraged a random forest, and SHAP values on top, to predict PICU admission. Overall, feedback from focus group sessions with clinicians indicated that the proposed solution was perceived as useful. However, the capacity to digest the information varied across clinical roles and levels of predictive modeling knowledge. In general, they preferred solutions that required less cognitive effort and could accommodate diverse user information needs. Still adopting ensemble methods, Kumar. et al. (2020) analyzed the agreement rate between physicians and LIME feature importances for sepsis prediction. Firstly, the authors trained several models such as random forest, adaboost, among others, and applied the LIME framework on top of the best-performing one (random forest). Secondly, physicians were asked to: (1) agree with the model whether the patient presented sepsis, based on individual features, and (2) blindly rank the top 3 features the model used to make predictions without accessing LIME explanations. Finally, LIME explanations were revealed and physicians were asked to rate their satisfaction on a 5-stars Likert scale. Results indicated that physicians agreed with model predictions (87%), and that LIME explanations were equal to the physicians’ explanations in at least two features for all patients (69%). Also, physicians were satisfied with LIME predictions (78%), and 68% presented a positive attitude towards explainability in terms of trust. However, such positive attitudes should not be translated into increased trust, as they found inconsistency in LIME explanations. Similarly in the scope of sepsis prediction, Sivaraman et al. (2023) examined clinicians’ interactions with an AI-based interpretable treatment recommender. The system was modeled using XGBoost, and SHAP as feature explainer. Through a think-aloud study involving 24 intensive care clinicians, the analysis revealed that explanations improved clinicians’ perceptions of the AI’s usefulness. In particular, the study identified four distinct behavior patterns among clinicians summarised as: Ignore, Negotiate, Consider, and Rely. Overall, clinicians who found the AI recommendations most useful saw it as additional evidence alongside their assessments.
@@
In the field of medical wearable technologies, Kim et al. (2022) developed MindScope, a system based on a personalized ML model trained on an XGBoost to predict patients’ stress levels. Interviews were conducted to gather feedback on the relevancy of feature importance (SHAP values) in three different explanation settings: (1) no explanations, (2) broad explanations, and (3) detailed explanations. Overall, it was found that the explanation-based settings were more useful than no explanations. However, detailed explanations were perceived as equally useful as broad explanations, thus questioning how detailed explanations should be.
@@
Next, Bienefeld et al. (2023) presented the development and evaluation of the Delayed Cerebral Ischemia Predictor (DCIP), a CDSS designed to forecast delayed cerebral ischemia in patients with aneurysmal subarachnoid hemorrhage. The system combined static and dynamic ML classifiers trained on multimodal patient data, with Extremely Randomized Trees selected as the best-performing model (AUC = 0.747). To enhance interpretability, SHAP values were integrated into the interface to display static and dynamic risk contributors. The study adopted a longitudinal multi-method evaluation with 112 participants, including an online survey with 95 ICU clinicians, a focus group, and 11 think-aloud interviews, analyzed through regression modeling and Grounded Theory. Results showed that while developers emphasized model interpretability, clinicians sought clinical plausibility, leading to design tensions around the utility of SHAP explanations. Overall, the study highlighted that bridging developer and clinician mental models requires ambidextrous XAI design strategies combining interpretability with clinically grounded plausibility.
@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
Next, in the field of oncology, Sabol et al. (2020) adopted a Cumulative Fuzzy Class Membership Criterion (CFCMC) to predict colorectal cancer. They developed two segmentation systems for whole-slide histopathological images: one using a standalone CNN and the other using the CFCMC classifier on CNN-extracted features to provide explanations. Through surveys with 14 pathologists, the CFMC showed higher acceptability compared to a plain CNN, pointing out that explainable classifiers improve AI usability and reliability in medical settings.
@@
Summary of Reviewed Studies
From our surveyed corpus (n = 17), the majority of systems used post-hoc explainability methods (14/17) while intrinsic interpretable models were less frequent (3/17). Model-agnostic approaches predominated (12/17) over model-specific methods (5/17). Outputs were roughly balanced between numerical and visual presentations (numerical: 10; visual: 7). The most frequently used XAI method was SHAP (5 studies), followed by LIME (2), integrated-gradient variants (IGrads, 2), with single occurrences of DeepTaylor, ICE/ALE, and model-specific saliency/variants; two studies reported transparent models without an added post-hoc XAI module. Human-centered evaluations were variable: Surveys (6), Think-aloud studies (3), Interviews (3), Mixed methods (3), and Focus groups (1). Notably, the majority of the studies (12/17) rely on a limited sample size to assess significance, i.e., lower than 30 participants, and lack of a structured protocol for HCE evaluation. Overall, these aggregated findings show a clear predominance of post-hoc, model-agnostic numerical explainers and a heterogeneous choice of HCE methods.
@@ 4 Challenges and Future Directions
4.1 Strengths & Limitations in XAI-based CDSS
@@
We have reviewed the diversity of CDSS that adopt XAI and human-centered evaluations in the medical field. Given the high-risk environment, application-grounded evaluations with domain experts were the main evaluation methodology. On average, clinical perceptions of XAI-based CDSS were mostly positive. For example, Ellenrieder et al. (2023) demonstrated that providing explanations not only enhances clinicians’ CDSS learning but also reduces false learning. In particular, Neves et al. (2021) demonstrated that less experienced practitioners could benefit from XAI explanations, while Brennan et al. (2019) revealed that adding XAI aids in significant improvements in physicians’ risk assessments.
@@
However, incorporating XAI still poses significant challenges and limitations. Firstly, Barda et al. (2020) discovered that the capacity to absorb explanations varied across clinical roles and levels of AI knowledge, thus advocating solutions that demand less cognitive effort. For the same reason, Kim et al. (2022) found that detailed explanations were perceived as equally useful as broad explanations, thus questioning how deep explanations should extend. Again, Kyrimi et al. (2020) documented that too-engineered explanations led to cognitive clutter, preventing their full internalization. Notwithstanding, Sivaraman et al. (2023) more granularly identified that clinicians who ignored explanations were already confident in their own settled decisions, but clinicians who accepted at least one aspect of the recommendation were able to reach a balanced intermediate solution that combined with their own intuition. Similarly, Matthiesen et al. (2021) argued that ML and explainability did not surpass specialized and experienced clinicians’ capabilities, but at best support them during the clinical workflow, as well as Kyrimi et al. (2020), who found that, although XAI was useful, current conventional clinical practices are still perceived as more understandable. Secondly, we highlight that only Hwang et al. (2022); Matthiesen et al. (2021) early gathered clinicians’ feedback before CDSS development, which may be crucial to ensure that the CDSS aligns with real-world clinical expectations, emphasizing the need for explanations to align with prior clinical knowledge through early stakeholder feedback to minimize the cognitive burden on clinicians. Thirdly, only few studies were conducted with reasonable sample sizes (>>30 participants: Bienefeld et al. (2023); Pumplun et al. (2023); Rajashekar et al. (2024); Ellenrieder et al. (2023); Kim et al. (2022)), limiting the ability to perform statistically significant inference, and lastly most CDSS-clinician interactions were made in simulated settings, lacking real workflow integration.
4.2 Impact of XAI on CDSS: a Socio-Technical Perspective
@@
Despite inherent challenges in explainability methods, such as lack of robustness Slack et al. (2020) and disagreement across methods and intra-method disagreement Krishna et al. (2022), issues that could be considered independent research topics per se, the friction from clinicians in perceiving explainability may be attributed to socio-technical gaps between human requirements and the effective utility of technological solutions, because human activity is “flexible and nuanced”, while computational mechanisms are “rigid and brittle” Ackerman (2000). In other words, a socio-technical gap refers to the discrepancy between human needs, including expectations, and the technological capabilities designed to support them, highlighting the limitations of technological systems in fully accommodating the nuances of human interactions. This general theoretical construct can be extended to the purpose of XAI explanations in CDSS, as explanation capabilities might not meet rigorous human requirements from precise clinical knowledge, given that CDSS are dramatically at high stake, as output recommendations may have a significant impact on patient lives.
@@
4.3 A Stakeholder-Centric Framework for CDSS Development
The development of CDSS requires an iterative, stakeholder-centric methodology that balances technical robustness with clinical usability. We propose a four-phase framework: (1) stakeholders and their requirements identification, (2) framing the research question and informing stakeholders, (3) prototype development, and (4) evaluation. Phases (3) and (4) are explicitly iterative, allowing for continuous development.
(1) Stakeholder Identification and Requirements
Core stakeholders include clinicians, patients, developers, and healthcare organizations (hospitals, clinics, management). However, for real-world adoption, regulators, payers/insurers, and policymakers must also be considered Shortliffe and Sepúlveda (2018). Stakeholder requirements are often divergent, as clinicians may emphasize usability, while managers may focus on efficiency, and regulators require compliance, necessitating to negotiate trade-offs van de Poel (2013).
Although our review shows that most empirical evaluations targeted clinicians, patients should also be considered as active stakeholders in XAI-based CDSS Ancker et al. (2021). Patients are directly affected by AI-informed decisions and increasingly encounter CDSS outputs in decision-making contexts. Explanations can play a key role in supporting informed consent, improving patient health literacy, and fostering trust in clinical recommendations. Yet, we found no systematic evaluations that directly assessed patient perceptions of XAI explanations. This absence highlights an important research gap: future work must design evaluation protocols that incorporate patient perspectives, ensuring that explanations are not only clinically useful but also understandable and actionable for patients.
(2) Framing the Research Question and Informing Stakeholders
Developers and researchers collaborate with stakeholders to define the clinical problem the CDSS will address. Shared problem framing ensures that system design aligns with actual decision-making needs rather than abstract performance benchmarks. Non-technical stakeholders must be informed of the CDSS’s underlying logic, including its machine learning models, XAI components, and evaluation frameworks. This stage establishes the evaluation criteria and metrics that guide subsequent development. To this end, Brankovic and others (2025) proposed a reference set of 14 evaluation metrics for XAI in CDSS divided into clinical attributes, decision attributes, and model attributes that all stakeholders should be aware of.
(3) Prototype Development
A progressive validation pipeline is recommended: initial testing with synthetic data, followed by retrospective studies on real-world datasets, and culminating in prospective trials Liu et al. (2019). Prototypes should be interoperable with EHR, respect privacy regulations, and integrate seamlessly into extant clinical workflows Weir et al. (2021). Human-centered usability testing, such as concurrent think-aloud studies, can capture clinicians’ reasoning and perceptions of system accuracy and explainability during interaction.
(4) Evaluation and Iteration
Evaluation serves both as a quality assurance mechanism and as feedback for redesign. It must address:
  • Technical validation: predictive accuracy, robustness, explainability, fairness, and safety.
  • Human-centered evaluation: trust, cognitive workload, workflow integration, and overall usability Khairat et al. (2018).
  • Ethical-legal considerations: accountability, transparency, and bias mitigation Ghassemi et al. (2021).
Phases (3) and (4) should be conceived as an iterative cycle, in which evaluation outcomes systematically inform redesign. This process should be guided by predefined checkpoints and measurable criteria (e.g., usability thresholds, calibration performance) to ensure that adjustments are targeted.
Field-Agnostic but Context-Sensitive
Although the framework is field-agnostic, it requires adaptation to local contexts. For instance, radiology-oriented CDSS may emphasize imaging pipelines, whereas psychiatry may require interpretability in narrative or unstructured data. Hence, the framework is best understood as generalizable but context-sensitive, consistent with socio-technical perspectives on CDSS design Greenes (2014).
5 Conclusion
@@
In this work, we examined the role of explainability in CDSS and the socio-technical challenges it faces. While explainability can enhance clinicians’ trust, its effectiveness varies with user expertise, cognitive load, and the depth of explanation required. The socio-technical gap remains a central challenge, as human decision-making is inherently flexible, whereas XAI models operate within rigid constraints. Given the high stakes in healthcare, rigorous evaluation strategies, particularly application-grounded methodologies, are essential to align CDSS with stakeholder needs. To address these challenges, we proposed an iterative framework that integrates XAI and HCI principles. By engaging key stakeholders and refining CDSS through continuous evaluation, we advocate for a development approach that prioritizes usability, trust, and clinical integration. In general, rather than replacing human expertise, CDSS should serve as augmentative tools that enhance clinicians’ decision-making. By synthesizing evidence on human-centered evaluations of XAI in CDSS and proposing a stakeholder-centric framework, this survey contributes to the advancement of trustworthy and integrative healthcare artificial intelligence, offering a roadmap for the future design and deployment of clinically viable CDSS.